OpenStudy (anonymous):

given f(x)=3x-7 what is f^-1(x)? A) 3x+7 B) -3x+7 C) (x+7)/3 D) (x/3)+7

OpenStudy (alexwee123):

inverse is switching the x and y around

OpenStudy (anonymous):

\[f(x) = 3x - 7\] \[f^-1(x) = 3y - 7\] \[x = 3y - 7\] \[x + 7= 3y\] \[(x + 7)/3= y\]

OpenStudy (anonymous):

thanks!